OpenStudy (anonymous):

what is the solution of this equation 3 + 1/4 X=2

OpenStudy (hockeychick23):

Ok first subtract 3 from both sides, what do you get?

OpenStudy (anonymous):

0 & 1

OpenStudy (anonymous):

http://sketchtoy.com/64167665

OpenStudy (hockeychick23):

not quite, 2-3= -1

OpenStudy (hockeychick23):

so once you subtract 3 from both sides your equation will look like this: 1/4x= -1. Do you know what to do next?

OpenStudy (anonymous):

no

OpenStudy (hockeychick23):

ok, so you want to get x by itself, so you need to multiply both sides of the equation by 4. So your answer is x= -1(4)

OpenStudy (anonymous):

thank you know how do i do it when it looks like this -1y= -4

OpenStudy (hockeychick23):

yes, you need to multiply both sides of the equation by -1 in that equation to get y by itself so its -1(-4)

OpenStudy (hockeychick23):

remember a negative*negative= a positive

OpenStudy (anonymous):

so the answer is 4 right ?

OpenStudy (hockeychick23):

correct! :)

OpenStudy (anonymous):

x+5=15 is x ten ?

OpenStudy (anonymous):

i mean id dint do anything just geussing

OpenStudy (hockeychick23):

yepp!

OpenStudy (anonymous):

no its 1.000.000

OpenStudy (hockeychick23):

you subtract 5 from both sides so its x=15-5 which is 10

OpenStudy (anonymous):

6y + 1 = -11 on which side do i mulitply

OpenStudy (hockeychick23):

you don't multiply, first subtract 1 from both sides, what do you get?

OpenStudy (anonymous):

-10

OpenStudy (hockeychick23):

not quite, its -11-1, so its like adding -1 + -11 which equals -12

OpenStudy (hockeychick23):

so your equation then looks like this: 6y= -12, next step is to divide both sides by 6, so what is -12/6

OpenStudy (anonymous):

sorry im little stupid lol

OpenStudy (hockeychick23):

no you're not(:

OpenStudy (anonymous):

-2

OpenStudy (hockeychick23):

correct so y= -2

OpenStudy (anonymous):

1/4 b = -3 will you help me lol theres only 3 more questions i just need somebody to go through them with me

OpenStudy (hockeychick23):

ok sure, so for this one you need to first multiply both sides of the equation by 4 to get b alone, so what is 1/4(4) ?

OpenStudy (anonymous):

1

OpenStudy (hockeychick23):

correct, now what is -3(4) ?

OpenStudy (anonymous):

-12

OpenStudy (hockeychick23):

correct, so b= -12

OpenStudy (anonymous):

-4x=-20

OpenStudy (hockeychick23):

ok for this one first divide both sides by -4, so what is -20/-4 ?

OpenStudy (anonymous):

5

OpenStudy (hockeychick23):

correct x=5

OpenStudy (anonymous):

1/2 z =10

OpenStudy (hockeychick23):

ok multiply both sides by 2, so 10(2)

OpenStudy (anonymous):

20

OpenStudy (hockeychick23):

correct z=20

OpenStudy (anonymous):

-7x-8=-36 Last one (: you helped me so much thank you i was working on this test for weeks and failed so many times

OpenStudy (hockeychick23):

no problem!(: ok so for this one first you need to add 8 to both sides, so what is -36+8? (its basically like doing 36-8 and adding a negative in front of it)

OpenStudy (anonymous):

-28

OpenStudy (hockeychick23):

correct so its -7x= -28, so what is -28/-7 ?

OpenStudy (anonymous):

4

OpenStudy (hockeychick23):

correct so x=4